(y+z)p+(z+x) =x+y


1
Expert's answer
2021-02-24T06:57:43-0500

"(y+z)p+(z+x)q =x+y"


"\frac{dx+dy+dz}{2(x+y+z)}=\frac{dx-dy}{y-x}"


"\frac{ln(x+y+z)}{2}=-ln(x-y)+lnc"


"(x+y+z)(x-y)^2=c_1"


"\frac{dx-dy}{y-x}=\frac{dy-dz}{z-y}"


"ln(x-y)=ln(y-z)+lnc_2"


"\frac{x-y}{y-z}=c_2"